# The Riemann Hypothesis is false

### v45

@@
$\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi\Big{(}\frac{1}{2}-\{{u\}}+O\Big{(}\frac{1}{1+||u||X}\Big{)}\Big{)}$ (30)
as $X\rightarrow\infty$. Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O\Big{(}\frac{1}{1+||T/2\pi||T^{1/4}}\Big{)}\Big{)}.$ (31)
Actually for $20<||u||=u<1/$ 2, the error term in (30) is $\ll\frac{1}{u(X+\frac{1}{2})}$ and for $||u||\geq 1/2$ it remains as it is 222see e.g. https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x. Thus, applying (31) into (30) with $u=T/2\pi m$ and splitting the error term for $20<||T/2\pi m||<1/$ 2 and $||T/2\pi m||\geq 1/2$ gives
@@
$+\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\Lambda(m)\Big{(}\frac{1}{2}+O(T^{-1/4})\Big{)}$ (35)
@@
$=(1-B+\log\pi)\frac{T}{2\pi}-\frac{T}{\pi}\log 2+o(T)$ (53)
for the sequence of $T,H\in[2T,2T+1]$ that satisfy (1). Suppose that $\Theta<1$. Then by Lemma 1 we have
@@
uniformly for $T\leq\gamma_{T}\leq H.$ Hence
@@
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{R}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
@@
Remark 3.
From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (49) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $(2n)^{it}$ by 1 for $2n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$. Note that the analogue of (49) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{(2n)^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite as elegant/useful as (49).
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need Theorem 4.11 of Titchmarsh [7] to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log t)^{-1}$. However, this is not possible (as can be seen from the approximate functional equation for $\zeta$). It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$. Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3). This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$. Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh assumes that $\sigma$ is bounded below by some constant $\sigma_{0}>0$.

### v46

@@
$\sum_{n\leq X}\frac{\sin(2\pi nu)}{n}=\pi\Big{(}\frac{1}{2}-\{{u\}}+O\Big{(}\frac{1}{1+X||u||}\Big{)}\Big{)}$ (30)
as $X\rightarrow\infty$. 222see e.g. https://math.stackexchange.com/questions/3331510/error-estimation-for-the-fourier-series-of-the-fractional-part-of-x Hence if $T\in\mathbb{Q}$ so that $T/2\pi m$ is irrational for $m\in\mathbb{N}$, we have
$\sum_{1\leq r\leq T^{1/4}}\frac{e^{\frac{irT}{m}}-e^{-\frac{irT}{m}}}{r}=\sum_{1\leq r\leq T^{1/4}}\frac{2i\sin(rT/m)}{r}=2\pi i\Big{(}\frac{1}{2}-\{{T/2\pi m\}}+O\Big{(}\frac{1}{1+T^{1/4}||T/2\pi m||}\Big{)}\Big{)}.$ (31)
Let $U_{m},V_{m}\in\mathbb{N},T=\frac{U_{m}}{V_{m}}$ and $m\mid U_{m}$. By Dirichlet’s theorem on rational approximations to irrationals, note that $\frac{1}{2\pi}$ can be approximated by infinitely many irreducible rational numbers $\frac{p_{m}}{q_{m}}$ with error $O({q_{m}}^{-2})$. Let: $V_{m}\mid p_{m}$, $\gcd(q_{m},U_{m})=1,\frac{U_{m}}{mV_{m}{q_{m}}^{2}}=o(1/q_{m})$ and $q_{m}\ll(\log T)^{6}$ uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Note that: $\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}$ is irreducible, $\Big{|}\Big{|}\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}\Big{|}\Big{|}\gg\frac{1}{q_{m}}$ and $\frac{T}{2\pi m}=\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}+O\Big{(}\frac{U_{m}}{mV_{m}{q_{m}}^{2}}\Big{)}=\frac{(\frac{U_{m}}{m})(\frac{p_{m}}{V_{m}})}{q_{m}}+o(1/q_{m})$. Hence $||T/2\pi m||\gg\frac{1}{q_{m}}\gg(\log T)^{-6}$ uniformly for $\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)$. Inserting (31) into (29) and applying the fact that $||T/2\pi m||\gg(\log T)^{-6}$, gives
@@
$+\sum_{\frac{T}{(1+\varepsilon)\log^{2}T}\leq m\leq T(1+\varepsilon)}\frac{\Lambda(m)}{2}$ (35)
@@
$=(1-B+\log\pi)\frac{T}{2\pi}+o(T)$ (53)
for the sequence of rational $T,H\in[2T,2T+1]$ that satisfy (1). Suppose that $\Theta<1$. Then by Lemma 1 we have
@@
uniformly for $T\leq\gamma_{T}\leq H.$ For large $x>0$, note that there are $\asymp x\log x$ zeros of $\zeta$ with imaginary parts in $(0,x]$ [5, Corollary 14.3]. Hence
@@
Remark 1. Notice that our argument actually shows that, for all large enough $T\in\mathbb{Q}^{+}$, there exists some $\rho_{T}$ with $T\leq\gamma_{T}\leq 2T+1$ and $\Re(\rho_{T})=1-o(1)$.
@@
Remark 3. From the literature on arXiv and elsewhere, it doesn’t appear like this approach on RH has been tried before. However, the idea of applying Gonek’s formula (49) in estimating sums over zeros of zeta in general, is not new. As one can confirm from the literature, it has proven to be a useful tool for such tasks. It is therefore instructive to sum (9) carefully, e.g. by not rushing to naively bound $(2n)^{it}$ by 1 for $2n$ not too close to $m$. Doing so will definitely make one lose important information needed for a sufficiently sharp bound for the sums over $2n\leq\frac{m}{1+\varepsilon}$ and $2n\geq m(1+\varepsilon)$. Note that the analogue of (49) from the Riemann explicit formula is $\sum_{T\leq\gamma_{T}\leq H}\frac{(2n)^{\rho_{T}-1}}{\rho_{T}}\ll\frac{(\log T)(\log\log T)}{T}.$ This has no main term, so it’s not quite as elegant/useful as (49).
Remark 4. The reader might wonder if our approach can give a lower bound for $\Theta_{T}:$ the supremum of the real parts of the zeros $\rho_{T}$ with $T\leq\gamma_{T}\leq H.$ For that, one would need our Lemma 1 to also hold for $\sigma=\sigma_{t}=o(1)$ with an error term $\ll(\log t)^{-1}$. However, this is not possible (as can be seen e.g by estimating $\zeta(\sigma+it)=\sum_{n\leq\frac{T}{2}}n^{-\sigma-it}=\sum_{n\leq T/2}n^{(-\sigma+\sigma_{t})-\sigma_{t}-it}$ for any fixed $\sigma>0$ via Abel summation, under the assumption that $\sum_{n\leq T/2}n^{-\sigma_{t}-it}=\zeta(\sigma_{t}+it)+O(1/\log t)$). It can also be seen that Titchmarsh’s proof of his Theorem 4.11 doesn’t hold for $\sigma=\sigma_{t}=o(1)$. Indeed, he begins his argument by quoting the Euler-McLaurin summation formula for $\zeta(\sigma+it)$ (equation 3.5.3). This is initially defined for $\sigma>1$, and then is extended by holomorphic continuation to $\sigma>0$. Recall that a function can only be holomorphically extended on open sets, and that’s precisely why Titchmarsh assumes that $\sigma$ is bounded below by some constant $\sigma_{0}>0$. For the same reason, one cannot apply any of the best known zero-free regions of $\zeta$ (e.g. Vinogradov-Koborov) to bound $\Re(\rho_{T})$ in (54).